OpenStudy (erinweeks):

Helllpp Please .. simplify the radical expression. -3√180h^2

OpenStudy (anonymous):

What this "h" means?

OpenStudy (erinweeks):

its just a variable

OpenStudy (erinweeks):

someone helpp pleasee !

OpenStudy (anonymous):

well that was wrong \[180=9\times 4\times 5\] so \[\sqrt{180}=\sqrt{9\times 4\times 5}=\sqrt{9}\times \sqrt{4}\times \sqrt{5}=3\times 2\times \sqrt{5}=6\sqrt{5}\]

OpenStudy (anonymous):

also \(\sqrt{h^2}=h\) so long as \(h\) is positive

OpenStudy (erinweeks):

but \[6\sqrt{5}\] isnt a answer for the quiz :/

OpenStudy (anonymous):

yes i see i ignored the \(-3\) out front not sure how the \(h\) becomes a \(p\) though

OpenStudy (anonymous):

you can finish it i am sure

OpenStudy (erinweeks):

nevermind i looked at the wrong question sorry ! its either A. \[6\sqrt{5h ^{4}}\] B. \[-18\sqrt{5h ^{4}}\] C. \[-18h ^{2}\sqrt{5}\] D. \[-3h \sqrt{90}\]

OpenStudy (anonymous):

maybe the original question is wrong, or maybe i read it incorrectly i read \[-3\sqrt{180h^2}\] and the answer is none of these

OpenStudy (anonymous):

Are you sure you posted the right question?

OpenStudy (anonymous):

question right*

OpenStudy (erinweeks):

ohh sorry guys im new to this it is \[-3\sqrt{180h ^{4}}\]

OpenStudy (anonymous):

it's okay.

OpenStudy (anonymous):

\[\sqrt{180h^4}= \sqrt{9\times4\times5 \times h^4}\] \[\sqrt4, \sqrt9 and \sqrt h^4\] are all perfect squares. Can you give me the answer for each one of them? :)

OpenStudy (anonymous):

\[\sqrt{h^4}\] *

OpenStudy (erinweeks):

\[\sqrt{4} = 2 \] \[\sqrt{9} = 3 \] i dont know what this would be \[\sqrt{h}^{4}\]

OpenStudy (anonymous):

oh i see now \(\sqrt{h^4}=\sqrt{(h^2)^2}=h^2\)

OpenStudy (erinweeks):

im confused

OpenStudy (anonymous):

easier to think : two goes in to four twice

OpenStudy (erinweeks):

is it one of the answers i posted though ?

OpenStudy (anonymous):

lets try it with numbers. lets put \(h=2\) so \(h^4=2^4=2\times 2\times 2\times 2=16\) then \[\sqrt{h^4}=\sqrt{16}=4=h^2\]

OpenStudy (anonymous):

lets work it out step by step \[-3\sqrt{180h ^{4}}\] \[=-3\sqrt{9}\sqrt{4}\sqrt{5}\sqrt{h^4}=-3\times 3\times 2\times h^2\times \sqrt{5}\] \[=-18h^2\sqrt{5}\]

OpenStudy (anonymous):

as usual, it is C it is always C

OpenStudy (erinweeks):

ok so you square them of first than multiiply them literally ?